SourMunchkin7806:

Give an example of a rational function that has a horizontal asymptote at y = 1 and a vertical asymptote at x = 4.

SourMunchkin7806:

@dude

dude:

So the same concept the denominator cannot equal 0 to get a vertical asymptote If \(\frac{1}{(x-1)}\) creates an asymptote at x=1, how would you get an asymptote at x=4?

SourMunchkin7806:

1 over (x-4)?

dude:

Vertical asymptote^

dude:

Right

SourMunchkin7806:

so would the answer be f(x)=1 over (x-4)?

dude:

Not quite, we are still looking for a horizontal asymptote at y=1

SourMunchkin7806:

so how do i get that?

dude:

Right now \(\frac{1}{x-4}\) Gives a horizontal asymptote at 0, do you have an idea on how to increase it by 1?

SourMunchkin7806:

ad one to the top?

dude:

Right now \(\frac{1}{x-4}\) Gives a horizontal asymptote at 0, do you have an idea on how to increase it by 1? The function is written as \(f(x)=\frac{a}{x-h}+k\) (Hint k translates up) Which is now \(f(x)=\frac{1}{x-4}+0\) Do you know how to translate it?

SourMunchkin7806:

No i dont

dude:

Add 1 to k x'D \(f(x)=\frac{1}{x-4}\color{green}{+1}\)

SourMunchkin7806:

ok awesome thanks lol

SourMunchkin7806:

i got one more can you help

dude:

Sure make a new post